# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ H.4 More Sensitivity Analysis
DeSGDA 77.6$\pm$0.9 84.3$\pm$1.1 70.5$\pm$0.6 84.8$\pm$1.4 76.6$\pm$0.7 83.9$\pm$0.9 71.9$\pm$0.6 76.9$\pm$1.1 76.1$\pm$0.8 73.7$\pm$1.0 76.0$\pm$1.2 70.4$\pm$0.7 76.9
[…]
WL subtree 73.5 79.5 64.8 75.9 58.9 68.4 72.5 72.0 69.7 63.6 76.1 74.0 70.7
GCN 51.2$\pm$0.1 71.1$\pm$0.4 42.7$\pm$0.4 27.8$\pm$0.3 32.1$\pm$1.1 27.0$\pm$0.2 55.2$\pm$0.6 50.5$\pm$0.7 50.9$\pm$1.1 49.1$\pm$0.3 67.1$\pm$0.6 57.3$\pm$0.6 48.5
GIN 66.9$\pm$2.2 78.9$\pm$2.3 60.3$\pm$3.1 72.8$\pm$0.3 51.1$\pm$0.6 68.6$\pm$1.8 63.5$\pm$2.1 67.8$\pm$3.7 65.9$\pm$1.7 60.3$\pm$1.8 71.1$\pm$1.1 67.2$\pm$1.3 66.2
GMT 50.9$\pm$0.5 73.0$\pm$0.1 57.3$\pm$0.3 73.0$\pm$0.4 66.5$\pm$0.2 73.0$\pm$0.3 72.4$\pm$0.6 50.9$\pm$0.1 66.5$\pm$0.4 58.3$\pm$0.2 66.5$\pm$0.5 72.8$\pm$0.3 65.1
CIN 60.1$\pm$0.7 73.1$\pm$1.1 57.5$\pm$0.2 73.0$\pm$0.4 66.5$\pm$1.1 73.1$\pm$0.7 58.5$\pm$2.1 52.9$\pm$1.4 66.5$\pm$1.3 56.1$\pm$0.1 66.5$\pm$0.4 57.4$\pm$0.7 63.4
SpikeGCN 63.3$\pm$0.4 72.6$\pm$0.7 60.6$\pm$0.2 73.1$\pm$0.8 65.4$\pm$0.6 66.6$\pm$1.2 64.3$\pm$1.4 64.7$\pm$1.0 65.6$\pm$1.2 59.8$\pm$0.8 70.1$\pm$1.7 60.9$\pm$1.4 65.6
DRSGNN 63.9$\pm$0.8 73.3$\pm$1.2 56.9$\pm$1.5 72.5$\pm$1.3 66.7$\pm$2.1 64.4$\pm$1.8 65.6$\pm$1.3 64.8$\pm$1.9 66.7$\pm$1.4 58.4$\pm$1.6 68.6$\pm$2.0 61.4$\pm$1.3 65.3
CDAN 57.1$\pm$0.4 75.0$\pm$0.7 61.2$\pm$0.4 73.7$\pm$0.1 68.2$\pm$0.4 73.3$\pm$0.3 60.2$\pm$0.1 56.5$\pm$1.4 68.2$\pm$0.2 53.9$\pm$1.4 68.4$\pm$0.2 59.6$\pm$0.5 64.6
ToAlign 49.1$\pm$0.3 27.0$\pm$0.6 57.3$\pm$0.5 27.0$\pm$0.4 66.5$\pm$0.5 27.0$\pm$0.2 57.3$\pm$0.3 49.1$\pm$0.4 66.5$\pm$0.2 49.1$\pm$0.3 66.5$\pm$0.1 57.3$\pm$0.4 50.0
MetaAlign 65.6$\pm$1.8 77.7$\pm$0.2 63.5$\pm$1.4 75.7$\pm$0.7 66.4$\pm$0.3 74.0$\pm$0.3 66.3$\pm$1.1 64.6$\pm$1.2 66.7$\pm$0.2 59.5$\pm$2.6 66.7$\pm$0.3 66.7$\pm$2.7 67.8
DEAL 64.0$\pm$0.9 71.9$\pm$1.2 61.4$\pm$0.3 73.3$\pm$0.3 64.9$\pm$1.4 71.9$\pm$1.9 62.5$\pm$2.1 66.2$\pm$0.5 54.2$\pm$1.4 55.6$\pm$0.8 64.6$\pm$0.4 58.8$\pm$0.4 64.1
CoCo 69.7$\pm$0.1 80.4$\pm$0.4 64.7$\pm$1.2 76.5$\pm$0.4 65.0$\pm$1.7 73.9$\pm$0.3 68.9$\pm$1.3 70.7$\pm$0.9 68.2$\pm$1.2 61.4$\pm$1.7 73.0$\pm$0.1 65.2$\pm$0.9 69.8
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 59.0$\pm$0.6 58.3$\pm$1.1 58.5$\pm$0.8 58.6$\pm$1.3 58.7$\pm$1.0 59.0$\pm$0.7 58.5$\pm$1.1 58.7$\pm$1.5 59.1$\pm$0.6 58.3$\pm$1.2 58.6$\pm$0.7 59.0$\pm$0.5 58.7
PA-BOTH 57.7$\pm$0.4 58.0$\pm$0.6 57.9$\pm$0.5 56.9$\pm$0.8 57.4$\pm$0.6 58.3$\pm$0.5 57.1$\pm$1.2 58.8$\pm$0.9 58.1$\pm$0.7 58.0$\pm$0.9 57.9$\pm$0.5 58.3$\pm$0.8 57.9
DeSGDA 64.4$\pm$1.2 76.9$\pm$1.5 64.9$\pm$0.9 76.6$\pm$1.2 68.6$\pm$1.8 74.1$\pm$1.3 66.9$\pm$0.8 65.1$\pm$1.2 69.9$\pm$1.5 63.9$\pm$2.0 70.9$\pm$1.6 64.2$\pm$1.4 68.9
[…]
WL subtree 65.7 71.8 57.9 71.1 47.4 43.4 65.5 75.1 45.3 34.9 52.7 49.8 56.7
GCN 70.6$\pm$2.1 60.3$\pm$1.5 60.5$\pm$3.4 62.3$\pm$1.1 58.4$\pm$0.5 43.2$\pm$0.2 63.8$\pm$1.0 70.3$\pm$0.3 50.6$\pm$1.0 32.8$\pm$0.3 50.1$\pm$0.4 42.2$\pm$0.2 55.4
GIN 66.7$\pm$2.1 73.7$\pm$2.4 57.3$\pm$3.1 69.4$\pm$2.3 58.6$\pm$0.4 43.1$\pm$0.3 66.4$\pm$2.7 74.8$\pm$1.8 42.2$\pm$1.6 33.5$\pm$1.0 57.4$\pm$0.8 43.9$\pm$2.3 57.2
GMT 67.3$\pm$0.3 56.8$\pm$0.4 58.0$\pm$0.2 56.8$\pm$0.2 60.6$\pm$0.3 56.8$\pm$0.5 57.8$\pm$0.1 67.3$\pm$0.1 39.5$\pm$0.3 67.3$\pm$0.2 39.5$\pm$0.5 57.8$\pm$0.4 57.1
CIN 67.6$\pm$0.4 63.7$\pm$2.1 58.9$\pm$1.0 56.8$\pm$0.4 63.6$\pm$0.4 59.5$\pm$2.7 58.7$\pm$1.2 67.0$\pm$0.5 61.7$\pm$1.6 67.8$\pm$0.7 62.2$\pm$2.1 56.0$\pm$1.3 61.9
SpikeGCN 67.2$\pm$0.5 57.2$\pm$1.2 57.9$\pm$0.8 57.1$\pm$0.6 61.1$\pm$1.3 58.9$\pm$1.6 60.0$\pm$1.2 67.2$\pm$0.9 53.9$\pm$2.1 64.4$\pm$0.8 57.8$\pm$1.0 59.9$\pm$1.2 60.2
DRSGNN 67.4$\pm$0.4 58.4$\pm$1.0 59.0$\pm$1.2 57.4$\pm$0.5 62.3$\pm$1.1 60.4$\pm$1.3 61.1$\pm$1.6 67.9$\pm$1.5 56.2$\pm$1.8 66.2$\pm$2.1 60.9$\pm$1.4 58.6$\pm$2.5 61.3
CDAN 72.9$\pm$0.4 72.7$\pm$0.4 65.4$\pm$0.3 72.9$\pm$0.1 61.2$\pm$0.3 70.3$\pm$0.2 65.7$\pm$0.4 72.7$\pm$0.1 61.0$\pm$0.1 72.1$\pm$1.2 60.7$\pm$0.2 65.3$\pm$0.6 67.7
ToAlign 32.7$\pm$2.0 43.2$\pm$0.1 42.2$\pm$1.3 43.2$\pm$0.9 60.5$\pm$0.7 43.2$\pm$1.2 42.2$\pm$0.4 32.7$\pm$1.2 60.5$\pm$0.9 32.7$\pm$0.3 60.5$\pm$0.7 42.2$\pm$0.4 44.7
MetaAlign 67.3$\pm$0.7 56.8$\pm$0.2 57.8$\pm$0.6 56.8$\pm$0.4 60.5$\pm$1.3 56.8$\pm$0.8 57.8$\pm$1.1 67.3$\pm$1.2 60.5$\pm$0.4 67.3$\pm$0.6 60.5$\pm$0.7 57.8$\pm$0.6 60.6
DEAL 75.0$\pm$0.9 76.3$\pm$2.4 65.9$\pm$1.8 77.5$\pm$2.7 60.3$\pm$4.5 69.7$\pm$3.2 67.2$\pm$1.5 75.3$\pm$1.7 57.4$\pm$4.1 71.1$\pm$2.2 65.7$\pm$2.7 66.4$\pm$1.6 69.0
CoCo 74.2$\pm$1.7 74.3$\pm$0.6 65.9$\pm$1.2 72.7$\pm$2.1 61.1$\pm$0.2 71.0$\pm$1.7 68.6$\pm$0.3 75.9$\pm$0.2 60.7$\pm$0.2 73.9$\pm$0.4 59.7$\pm$1.1 67.3$\pm$0.8 68.8
SGDA 55.9$\pm$0.6 57.1$\pm$0.5 56.1$\pm$0.4 54.6$\pm$0.8 55.8$\pm$1.1 57.7$\pm$0.6 54.3$\pm$0.7 53.6$\pm$1.3 59.1$\pm$0.8 56.7$\pm$0.6 55.4$\pm$1.2 53.8$\pm$0.5 55.9
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.9$\pm$0.7 55.7$\pm$0.4 56.6$\pm$0.6 57.1$\pm$1.0 56.1$\pm$1.2 55.8$\pm$0.5 56.5$\pm$0.7 55.5$\pm$0.4 55.9$\pm$0.8 56.2$\pm$0.6 56.5$\pm$1.5 56.0$\pm$0.5 56.2
PA-BOTH 56.4$\pm$0.5 55.9$\pm$0.6 56.0$\pm$0.5 56.4$\pm$0.4 56.3$\pm$0.6 57.7$\pm$0.7 56.6$\pm$0.2 58.8$\pm$0.9 56.9$\pm$0.7 57.2$\pm$0.3 56.5$\pm$0.5 58.3$\pm$0.8 56.9
DeSGDA 75.2$\pm$0.8 74.3$\pm$1.2 68.0$\pm$1.5 73.6$\pm$2.0 62.3$\pm$1.7 71.2$\pm$2.5 68.8$\pm$1.3 76.0$\pm$1.2 61.9$\pm$2.2 71.4$\pm$1.6 62.8$\pm$1.2 65.7$\pm$0.9 69.3
[…]
WL subtree 78.0 68.7 70.1 70.5 59.0 61.2 71.7 78.0 49.9 56.3 69.4 71.9 67.1
GCN 74.5$\pm$0.2 60.8$\pm$2.1 69.7$\pm$0.4 68.5$\pm$1.7 54.1$\pm$0.9 55.2$\pm$0.9 68.6$\pm$1.6 75.5$\pm$0.5 51.5$\pm$1.3 46.4$\pm$1.7 58.6$\pm$0.4 60.2$\pm$0.2 61.9
GIN 77.9$\pm$3.1 70.7$\pm$2.4 70.9$\pm$0.8 69.2$\pm$1.2 64.1$\pm$1.0 61.9$\pm$2.4 78.5$\pm$0.2 79.8$\pm$3.3 65.5$\pm$2.7 71.5$\pm$0.9 69.5$\pm$1.8 73.5$\pm$2.6 71.1
GMT 67.3$\pm$0.2 52.5$\pm$0.1 59.9$\pm$0.3 47.5$\pm$0.2 53.5$\pm$0.2 52.5$\pm$0.4 59.9$\pm$0.1 67.3$\pm$0.2 46.7$\pm$0.5 67.3$\pm$0.3 53.3$\pm$0.1 59.9$\pm$0.4 57.1
CIN 70.8$\pm$1.1 66.9$\pm$3.4 61.7$\pm$0.6 62.6$\pm$2.4 56.3$\pm$3.1 62.9$\pm$1.3 65.1$\pm$1.0 68.8$\pm$1.7 56.6$\pm$1.4 66.9$\pm$1.0 58.1$\pm$1.3 62.5$\pm$0.9 63.3
SpikeGCN 63.2$\pm$0.6 55.9$\pm$1.7 59.8$\pm$1.2 56.5$\pm$1.6 54.1$\pm$2.2 60.3$\pm$1.2 60.1$\pm$0.8 68.8$\pm$1.4 54.4$\pm$2.0 64.2$\pm$1.6 56.6$\pm$0.8 58.5$\pm$1.5 59.4
DRSGNN 56.9$\pm$0.9 53.4$\pm$1.2 58.9$\pm$1.0 57.7$\pm$1.4 53.2$\pm$1.6 61.2$\pm$1.5 60.6$\pm$2.1 67.9$\pm$1.2 57.1$\pm$1.4 67.3$\pm$2.0 59.9$\pm$1.9 57.1$\pm$0.8 59.3
CDAN 75.5$\pm$0.1 71.3$\pm$0.4 70.7$\pm$0.3 70.3$\pm$0.1 58.7$\pm$0.6 58.4$\pm$0.6 70.2$\pm$0.5 76.1$\pm$0.5 58.5$\pm$0.6 69.4$\pm$1.5 59.0$\pm$0.1 63.7$\pm$1.4 66.8
ToAlign 67.3$\pm$0.2 47.5$\pm$0.4 59.9$\pm$0.6 47.5$\pm$0.5 46.7$\pm$0.4 47.5$\pm$0.2 59.9$\pm$0.7 67.3$\pm$0.3 46.7$\pm$0.1 67.3$\pm$0.4 46.7$\pm$0.5 59.9$\pm$0.3 55.4
MetaAlign 76.5$\pm$0.4 71.8$\pm$1.1 71.8$\pm$0.8 71.4$\pm$0.9 59.3$\pm$0.8 63.0$\pm$1.0 74.2$\pm$1.6 78.0$\pm$0.2 61.7$\pm$1.2 69.9$\pm$1.6 62.2$\pm$0.4 68.3$\pm$1.5 69.0
DEAL 76.6$\pm$0.8 68.8$\pm$1.0 69.9$\pm$0.4 66.4$\pm$0.8 59.3$\pm$2.1 64.2$\pm$2.2 79.1$\pm$0.1 81.9$\pm$0.6 64.5$\pm$1.1 75.3$\pm$0.6 69.8$\pm$1.6 76.5$\pm$0.2 71.0
CoCo 75.5$\pm$0.4 71.7$\pm$0.7 68.7$\pm$1.1 69.2$\pm$2.0 60.8$\pm$1.1 65.7$\pm$0.3 79.2$\pm$1.2 76.8$\pm$0.6 63.8$\pm$0.5 73.8$\pm$0.4 64.6$\pm$0.8 70.1$\pm$1.1 70.0
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.4$\pm$0.6 56.3$\pm$0.2 55.6$\pm$0.8 55.1$\pm$0.5 55.3$\pm$1.1 55.9$\pm$0.4 56.1$\pm$0.7 55.7$\pm$0.6 57.1$\pm$0.3 56.6$\pm$1.2 55.2$\pm$0.7 56.8$\pm$1.0 55.9
PA-BOTH 55.9$\pm$1.0 56.0$\pm$0.5 56.1$\pm$0.7 56.6$\pm$1.2 55.9$\pm$0.6 56.0$\pm$0.7 57.3$\pm$0.8 56.8$\pm$1.3 55.9$\pm$1.2 56.3$\pm$1.0 56.4$\pm$0.9 57.1$\pm$1.3 56.4
DeSGDA 76.7$\pm$1.0 72.5$\pm$0.6 72.0$\pm$2.3 71.5$\pm$1.5 61.0$\pm$1.2 67.6$\pm$0.8 75.5$\pm$1.0 79.8$\pm$1.2 62.8$\pm$1.6 75.9$\pm$1.3 65.5$\pm$2.1 72.3$\pm$1.7 71.1
@@
WL subtree 72.6 80.3 62.7 75.5 52.0 63.6 69.1 69.8 70.7 59.4 80.0 70.6 68.9
GCN 49.5$\pm$0.4 71.1$\pm$0.4 46.8$\pm$0.5 33.7$\pm$2.8 32.7$\pm$0.4 27.4$\pm$0.1 56.2$\pm$1.5 55.3$\pm$0.4 58.2$\pm$1.7 51.0$\pm$0.2 60.7$\pm$3.7 53.2$\pm$0.2 49.6
GIN 67.3$\pm$2.7 67.9$\pm$4.8 61.5$\pm$4.2 65.4$\pm$3.7 58.9$\pm$4.1 61.0$\pm$3.4 62.5$\pm$3.2 66.2$\pm$2.1 69.7$\pm$0.9 56.8$\pm$0.7 72.4$\pm$2.8 64.0$\pm$1.6 64.5
GMT 50.3$\pm$1.2 42.5$\pm$3.4 51.1$\pm$3.7 42.5$\pm$4.5 56.1$\pm$4.7 42.5$\pm$4.1 53.2$\pm$4.9 51.0$\pm$0.2 68.2$\pm$0.4 51.0$\pm$0.3 68.2$\pm$0.5 53.2$\pm$0.4 52.5
CIN 51.1$\pm$0.2 72.6$\pm$0.1 54.0$\pm$0.9 72.6$\pm$0.2 68.2$\pm$0.3 71.5$\pm$1.3 55.0$\pm$2.1 53.5$\pm$1.8 68.2$\pm$0.3 52.0$\pm$0.3 68.3$\pm$0.1 53.6$\pm$0.6 61.7
SpikeGCN 62.8$\pm$0.8 73.1$\pm$1.2 61.4$\pm$0.8 70.9$\pm$0.6 57.7$\pm$1.6 66.2$\pm$1.1 61.2$\pm$1.6 64.5$\pm$1.0 62.3$\pm$1.4 57.3$\pm$0.9 68.9$\pm$1.2 60.1$\pm$1.0 63.6
DRSGNN 64.3$\pm$0.6 76.3$\pm$0.9 56.7$\pm$1.1 73.2$\pm$0.8 58.6$\pm$1.4 63.9$\pm$1.8 63.0$\pm$2.1 65.1$\pm$1.6 64.1$\pm$1.9 59.2$\pm$2.2 70.8$\pm$2.5 56.6$\pm$1.4 64.3
CDAN 59.6$\pm$0.3 73.8$\pm$0.5 56.7$\pm$1.4 73.7$\pm$0.3 71.2$\pm$0.4 73.2$\pm$0.3 55.5$\pm$0.2 57.3$\pm$1.1 69.9$\pm$0.2 54.6$\pm$2.0 69.8$\pm$1.4 56.6$\pm$0.3 64.3
ToAlign 51.0$\pm$0.2 27.4$\pm$0.1 53.2$\pm$0.4 27.4$\pm$0.2 68.2$\pm$0.3 27.4$\pm$0.3 53.2$\pm$0.1 51.0$\pm$0.2 68.2$\pm$0.2 51.0$\pm$0.4 68.2$\pm$0.3 53.2$\pm$0.2 50.0
MetaAlign 65.0$\pm$0.7 77.6$\pm$1.6 62.0$\pm$0.6 77.1$\pm$0.9 68.2$\pm$0.8 74.5$\pm$2.0 64.2$\pm$0.9 65.4$\pm$0.3 68.0$\pm$0.3 56.1$\pm$2.3 68.2$\pm$0.1 66.2$\pm$1.1 67.7
DEAL 65.6$\pm$0.6 73.0$\pm$0.9 58.0$\pm$0.3 71.6$\pm$1.6 60.1$\pm$2.8 73.1$\pm$0.5 62.8$\pm$1.0 65.0$\pm$2.4 65.8$\pm$0.8 53.9$\pm$2.6 57.6$\pm$2.8 56.7$\pm$3.1 63.6
CoCo 70.4$\pm$0.7 80.4$\pm$0.9 62.4$\pm$0.8 75.8$\pm$1.2 65.7$\pm$2.0 73.7$\pm$0.3 67.0$\pm$0.8 70.4$\pm$0.7 69.7$\pm$0.4 62.7$\pm$0.9 74.4$\pm$0.5 63.7$\pm$0.9 69.7
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 59.2$\pm$0.8 58.7$\pm$0.5 59.0$\pm$1.1 58.7$\pm$0.8 58.9$\pm$0.6 59.2$\pm$1.2 58.7$\pm$0.6 58.6$\pm$1.2 59.0$\pm$1.0 59.5$\pm$0.6 58.7$\pm$0.5 58.5$\pm$1.1 58.9
PA-BOTH 57.6$\pm$0.5 58.4$\pm$0.4 58.9$\pm$0.6 57.4$\pm$0.6 57.1$\pm$1.0 58.4$\pm$0.5 58.0$\pm$1.0 58.1$\pm$0.5 58.4$\pm$0.6 57.7$\pm$1.1 57.5$\pm$0.6 58.0$\pm$0.4 58.0
DeSGDA 66.0$\pm$0.5 76.2$\pm$0.4 62.8$\pm$0.6 77.6$\pm$0.6 68.5$\pm$1.0 74.6$\pm$0.5 65.1$\pm$1.0 65.4$\pm$0.5 70.9$\pm$0.6 64.8$\pm$1.1 72.6$\pm$0.6 66.3$\pm$0.4 69.3

### v4

@@ Appendix I Baselines
DeSGraDA 76.3$\pm$1.9 84.6$\pm$2.5 69.2$\pm$2.3 83.6$\pm$2.6 77.5$\pm$2.2 83.7$\pm$1.9 69.8$\pm$2.4 74.0$\pm$1.6 76.2$\pm$2.0 73.0$\pm$2.1 77.8$\pm$2.3 70.5$\pm$1.7 76.4
@@
WL subtree 49.2 56.8 29.6 20.1 21.0 18.4 59.5 50.5 57.3 48.1 63.9 66.9 49.3
GCN 48.9$\pm$2.8 59.0$\pm$1.7 20.7$\pm$2.0 27.3$\pm$2.3 15.1$\pm$1.8 26.9$\pm$2.2 61.6$\pm$1.9 53.6$\pm$1.5 68.1$\pm$1.6 52.9$\pm$2.6 64.9$\pm$2.1 69.7$\pm$2.3 47.4
GIN 48.8$\pm$1.9 24.7$\pm$2.1 44.1$\pm$1.8 22.4$\pm$2.3 57.0$\pm$2.1 18.4$\pm$2.0 73.0$\pm$1.8 52.5$\pm$2.3 63.2$\pm$1.6 53.6$\pm$1.5 70.3$\pm$2.6 69.4$\pm$1.8 49.6
GMT 49.1$\pm$1.9 32.9$\pm$2.2 31.8$\pm$1.8 27.3$\pm$2.3 52.5$\pm$1.5 27.6$\pm$1.8 75.4$\pm$1.9 53.2$\pm$2.1 74.1$\pm$2.6 57.9$\pm$2.4 70.9$\pm$1.8 71.1$\pm$2.7 52.0
CIN 50.4$\pm$1.8 18.4$\pm$2.0 21.2$\pm$2.1 36.8$\pm$1.8 43.0$\pm$2.1 22.9$\pm$1.9 53.4$\pm$1.7 56.5$\pm$1.5 62.3$\pm$1.6 53.3$\pm$1.9 75.0$\pm$2.1 69.3$\pm$2.0 51.4
SpikeGCN 53.4$\pm$1.3 61.5$\pm$1.1 28.4$\pm$2.1 48.0$\pm$1.4 20.8$\pm$1.4 51.2$\pm$1.4 69.3$\pm$2.0 62.4$\pm$1.6 77.9$\pm$1.1 64.1$\pm$1.5 76.8$\pm$2.1 71.9$\pm$1.8 57.1
DRSGNN 52.1$\pm$1.5 69.7$\pm$1.8 28.7$\pm$1.7 42.4$\pm$2.1 18.6$\pm$2.0 48.3$\pm$2.8 79.5$\pm$2.0 59.5$\pm$1.5 77.7$\pm$1.9 63.1$\pm$1.5 76.4$\pm$2.1 72.1$\pm$2.3 57.4
CDAN 49.6$\pm$2.2 69.8$\pm$1.6 44.1$\pm$1.8 33.9$\pm$1.9 43.1$\pm$2.4 42.3$\pm$2.0 70.5$\pm$1.8 60.3$\pm$2.2 76.6$\pm$2.1 60.1$\pm$1.4 75.8$\pm$2.5 70.5$\pm$2.3 58.1
ToAlign 54.0$\pm$2.4 71.0$\pm$2.1 34.8$\pm$1.8 46.9$\pm$1.9 29.6$\pm$2.3 45.7$\pm$1.8 71.9$\pm$2.2 61.6$\pm$2.1 76.7$\pm$1.9 62.8$\pm$2.3 76.4$\pm$2.0 71.0$\pm$1.3 58.5
MetaAlign 48.1$\pm$2.0 70.0$\pm$1.9 30.7$\pm$1.4 18.4$\pm$1.8 24.9$\pm$2.3 18.4$\pm$1.9 70.1$\pm$2.3 51.9$\pm$1.5 74.6$\pm$2.4 51.9$\pm$2.2 75.1$\pm$1.8 69.3$\pm$1.7 52.7
DEAL 57.9$\pm$2.3 71.6$\pm$2.0 57.2$\pm$1.8 59.3$\pm$1.5 62.2$\pm$2.2 59.4$\pm$1.9 72.1$\pm$2.3 63.9$\pm$1.7 78.2$\pm$2.2 62.6$\pm$1.8 78.3$\pm$2.1 77.3$\pm$1.9 66.5
CoCo 59.5$\pm$2.1 70.4$\pm$1.7 56.6$\pm$2.6 58.3$\pm$2.2 59.4$\pm$1.9 53.9$\pm$2.5 74.7$\pm$1.4 62.7$\pm$2.0 70.6$\pm$1.6 63.1$\pm$2.0 77.2$\pm$1.7 76.4$\pm$2.4 65.2
SGDA 57.7$\pm$1.5 63.8$\pm$2.1 49.8$\pm$2.3 54.1$\pm$1.7 42.6$\pm$2.3 54.9$\pm$1.6 74.1$\pm$2.8 63.0$\pm$2.3 78.7$\pm$2.7 64.5$\pm$2.1 76.8$\pm$2.2 74.2$\pm$1.6 62.9
StruRW 50.0$\pm$2.3 53.1$\pm$1.9 32.4$\pm$2.4 40.6$\pm$2.1 26.0$\pm$2.4 38.4$\pm$2.0 73.3$\pm$1.6 61.7$\pm$1.8 71.2$\pm$1.9 53.6$\pm$2.3 75.2$\pm$2.1 71.0$\pm$2.2 53.9
A2GNN 56.1$\pm$2.0 68.5$\pm$1.6 48.7$\pm$2.1 52.5$\pm$1.8 42.9$\pm$1.4 48.4$\pm$1.8 70.8$\pm$1.7 51.9$\pm$2.0 76.3$\pm$2.2 51.9$\pm$1.8 75.1$\pm$1.6 69.3$\pm$2.4 59.4
PA-BOTH 51.4$\pm$1.8 62.7$\pm$2.1 31.8$\pm$2.0 40.5$\pm$2.3 28.5$\pm$1.9 45.0$\pm$2.2 69.5$\pm$1.6 61.0$\pm$1.5 68.4$\pm$1.7 57.8$\pm$2.2 73.0$\pm$2.4 73.3$\pm$2.5 55.3
DeSGraDA 60.1$\pm$2.2 76.1$\pm$1.8 63.9$\pm$1.9 71.7$\pm$2.0 61.1$\pm$1.9 62.3$\pm$1.6 73.6$\pm$2.3 64.2$\pm$2.0 80.9$\pm$2.2 64.7$\pm$1.8 82.0$\pm$2.5 80.6$\pm$2.1 70.1